# The SABRE South Technical Design Report Executive Summary

### v2

@@ 4.2 R11065 PMTs
e R11065 PMTs were chosen for their high sensitivity to single photons, low noise rates, and low radioactive background. Only PMTs with quantum efficiency (QE) $>30\%$ were procured, so to achieve a high single-photon detection efficiency. Custom PMT voltage dividers are used to minimise their contribution to the background. These bases will follow the resistance ratios recommended by Hamamatsu for this type of PMT. SABRE South has decided on a design that will use a positively biased voltage divider. The circuit diagram and an image of the PCB for the voltage divider design are shown in Fig. 7. The final design is expected to use small ceramic capacitors.
Figure 7: Left: The circuit diagram for the final R110065 base design. Right: One of the PCBs to be used for the base circuit, which will be attached to an R11065 PMT.
The pre-calibration program for the R11065 (crystal detector) PMTs focuses on confirming their single-photon detection efficiency, characterising their single-photo electron (PE) PMT response, and characterising the noise as a function of temperature and voltage. These parameters are used as an early calibration input to SABRE South data analysis and as an input into the digitisation stage of detector simulations. The requirements for a quantum efficiency larger than $30\%$ (at a wavelength of 420 nm) and excellent signal-to-background discrimination for single photons are driven by the light yield requirements (of $\sim$ 12-14 PE/keV) on the crystal detectors to achieve a threshold of 1-2 keV. Several essential measurements will be completed for all PMTs under strict cleanliness protocols. These are: relative quantum efficiency, gain, dark rate, and afterpulse rate at nominal voltage. Further characterisation of single-photon response and noise, including temperature and bias voltage dependence, is performed on a limited set of PMTs to limit overall exposure. As SABRE operates at room temperature, dark counts due to thermionic emission and temperature dependence of this dark rate are essential factors that are more easily measured in a dedicated setup than in-situ. The precalibration datasets allow for the development of data analysis techniques for discrimination of thermionic emission from low-energy events in the crystals, near the 1-2 keV region. There have previously been reports [19, 20, 21] of light emission likely due to anode and dynode glow following a sufficiently intense light signal in the R11065 and the related R11410 PMT types. Such light emission could introduce an additional background in SABRE and therefore needs to be characterised if present.
@@ 4.3 Cleanliness protocols
All components of the crystal enclosures are cleaned and vacuum baked prior to final assembly to ensure that there is no moisture remaining on the components. The crystal PMTs, cables, and bases are cleaned with lint-free isopropyl alcohol (IPA) soaked cleanroom wipes and a high pressure nitrogen gun to remove any debris. To prevent radon plate-out, all crystal detector module components are protected in two-layered radon bags in-between cleaning and their installation in the final detector modules. The assembly of the base/cables and their mounting on the PMTs is performed in a cleanroom environment. Assembly of the crystal modules is performed in a nitrogen flushed glove box.
@@ 4.5 Gas Handling System
Figure 10: The exhaust system gas panel undergoing laboratory testing at ANU.
@@ 5 Liquid scintillator veto
SABRE South’s active veto system serves to identify background events originating from radioactive decays within the crystals that may mimic a signal and to act as passive shielding for background external to the crystals. For example, 40K, a key impurity, decays via the emission of an Auger electron with an energy in the range of 1-6 keV coincdent with a 1.5 MeV $\gamma$ ray. Previous experiments have accounted for this by employing double-coincidence rejection, where it was assumed that the escaping gamma ray would be detected by another crystal in the sensitive mass [22]. However, this detection mechanism is not fully efficient, as demonstrated by the peak at around 3 keV in the DAMA data where events from 40K are expected [23]. The key components of the LS veto detector are summarised in Table 4.
By immersing the NaI(Tl) crystal detector modules in a liquid scintillator (Fig. 11), high-energy $\gamma$ rays produced by radioactive decays can be detected and the decay event vetoed. The SABRE South full background simulation demonstrated that the veto can reduce background from important 40K contamination by a factor of 10 in the region of interest, and suppresses the total experimental background by 27% [16].
@@
SABRE South uses approximately 11,600 L of liquid scintillator that is a mixture of linear alkylbenzene (LAB), sourced from the JUNO experiment, doped with approximately 3.5 g/L of the fluorophore PPO (2,5-diphenyloxazole) and 15 mg/L of bis-MSB (1,4-bis(2-methylstyryl) benzene). The fluorophore concentrations are subject to further optimisation to ensure low absorption. The veto detector uses Hamamatsu 204 mm R5912 PMTs with oil-proof potted bases. The nominal design uses a baseline of 18 PMTs, but it is possible a total of 32 could be installed with the inclusion of PMTs from the decommissioned Daya Bay experiment. The PPO and Bis-MSB are required to mitigate the self-absorption properties of LAB [24] and maximise the total light yield. The wavelength of light output of LAB alone is around 300 nm [25]. In the presence of fluorophores, however, the maximum wavelength of emission is 422 nm, which is closer to the peak quantum efficiency of the PMTs.
@@
The liquid scintillator is effective in rejecting events in which electrons, neutrons, and photons deposit more than 50 keV of energy in the scintillator. For the baseline scenario with 18 veto PMTs, this enables the detection of energy depositions with 50% efficiency at a threshold of 50 keV, rising to 99% at 200 keV. This expected performance is important if the DAMA/LIBRA signal is to be excluded, as it provides a means of understanding what background processes might cause a DAMA/LIBRA-like modulation through use of techniques such as pulse-shape discrimination. This has been tested conceptually using an LS prototype detector, containing approximately 40 L of LAB-based LS and a single R5912 PMT, in which neutron and $\gamma$ ray signals could be discriminated. This suggests that the same approach could be applied to the full LS veto, particularly since the arrival of 16 additional PMTs from the decommissioned Daya Bay experiment gives us the opportunity to increase the number of PMTs in the detector from 18 to 32, thereby increasing photosensor coverage and light yield/light collection efficiency.
@@ 5.1 Vessel design
The vessel structure can be divided into three sections: the central barrel and the two torispherical caps located at the top and bottom of the vessel. To ensure that the vessel is small enough to be safely transported to SUPL on the back of a vehicle and large enough for MeV-scale photons to be fully absorbed inside the LS, an outer diameter of 2600 mm and total height of 3322 mm were chosen. A diagram of the vessel is shown in Fig. 12. Rows of stud bolts are welded to the internal wall of the vessel to mount the PMTs and reflector. A central flange (“top flange”, see Fig. 13), is located at the top of the vessel, which is flanked by a circle of smaller side flanges. The central flanges are used for the insertion of the crystal enclosures, fluid handling, cable feedthroughs, and calibration systems.
@@ 5.2 Assembly
Figure 14: Engineering models of the central and lateral mounting units. The mounts are nominally made of aluminium. Labels 1-3 highlight the central mount frame; labels 4 and 5 point out the bulb and oilproof base of the R5912 respectively; and labels 6-8 indicate the clamps in which the PMT is mounted. The mounts will be fastened to two stud bolts, at the top and bottom of each mounting bracket, whilst the clamps will be fastened with stainless steel nuts and bolts.
@@ 5.3 Cleanliness protocols
The vessel must be pickled internally because an oxide layer is expected to have formed between manufacturing and assembly, which may contain radioactivity and may not be compatible with the LS. Pickling is performed using a corrosive solution with a concentration that is balanced enough to corrode the oxide layer and not the steel itself. Following this, the vessel is rinsed multiple times with a solution of de-ionised water and Alconox detergent, followed by a rinse with de-ionised water. Rinsing ceases once the outgoing water/solution reaches a required resistivity of $\geq 14$ MOhm cm. The R5912 veto PMTs and Lumirror follow a similar cleaning procedure to the R11065 PMTs, and also performed in a cleanroom environment. The installation of internal components is performed in a clean room-like environment with radon reduction, with all installation personnel wearing clean-suits, boot-covers, gloves, and hairnets.
@@ 5.4 R5912 PMTs
Owing to the sparse array of PMTs relative to the volume of LS, the probability of photoelectron (PE) detection is 0.2-0.3 PE/keV/PMT. Consequently, the pre-calibration of these PMTs is driven by the need to be sensitive to low numbers of PEs, to inform our simulation/digitisation and thus analysis, and also by the ability to deploy a degree of position reconstruction and particle ID in the LS veto. With such reconstruction capabilities, we can disentangle our background contributions and pursue other physics using the veto sub-detector — for example, studies of neutron spallation induced by muon tracks, or potential detection of supernova neutrinos (in which tagging the neutron of an inverse beta decay could improve our detection efficiency).
@@ 6 Muon veto
The digitiser used for the readout is a CAEN V1743, which was specifically chosen for its high sample rate of 3.2×$\times 10^{9}$ samples/s (which matches the TTS of the PMTs) and resolution of $<$ 8 ps RMS [28]. The timing resolution of this system allows for accurate position reconstruction along the length of each muon paddle, where the time difference between signals in each PMT is used to determine position. Measurements performed on the surface have found that this can be performed at a position resolution of approximately 5 cm. The efficiencies of each detector have also been measured on the surface to be at least 99% for muons that fully traverse the thickness of the detector. The surface flux at the University of Melbourne was measured to be 131.3 $\pm$ 0.4 s-1m-2, while at SUPL we expect the muon rate for one panel (1.2 m2) to be 37.8 $\pm$ 4.3 d-1, and 302.7 $\pm$ 34.0 d-1 for all eight panels with an overall coverage of 9.6 m2.
@@
Figure 16: Muon detectors in SUPL, assembled in ‘telescope’ mode, to measure the muon flux and angular spectrum.
@@ 7 Calibration systems
The SABRE South sub-detectors require periodic calibration, which is particularly important for accurate time-dependent measurements. Calibration refers to the process of bias-correcting and verifying the accuracy of a detector’s measurements. The LS veto and NaI(Tl) calibration systems for SABRE South are designed for remote operation, and are located within the shielding structure. The response of the detector to particle interactions can change with time, due to: (i) environmental changes such as temperature, humidity, mechanical stress and vibrations (recorded by the slow control system outlined in Sec. 8.5); (ii) electrical changes: variations in the supply voltage, EM interference; (iii) detector ageing: over time optical properties of scintillators can deteriorate (e.g. reduced absorption length or light yield), and the PMTs can vary in detection efficiency and other features used for pulse shape discrimination.
SABRE South uses radioactive sources for the calibration of the NaI(Tl) detectors, the liquid scintillator veto, and the muon detector. The liquid scintillator veto is also calibrated with an optical (pulsed laser light) system. The two liquid scintillator veto calibration systems are complementary, with the radioactive calibration system providing accurate high-energy calibration lines and the optical system capable of sending low-intensity high-frequency pulses to test the single photoelectron and timing response of the PMTs. Calibration with radioactive sources requires data runs of a few hours every time a calibration run is scheduled (other NaI(Tl) experiments have 3–4 hour calibration runs every 2 weeks [29]). The laser calibration can be run more frequently in short bursts as it does not require any mechanical adjustments to the system. The frequency of calibration of the two systems will be adjusted as necessary to optimise for detector stability and live-time.
@@ 7.1 Radioisotope calibration system
Radioactive sources must be placed near the detectors in order to calibrate their responses. During these periods (calibration runs), the signal induced by the calibration source must be strong enough to be distinguishable from the other detected signals. Conversely, during normal data acquisition, the signal induced by the calibration source must be negligible. This is achieved by keeping the sources shielded and at a distance from the detectors during normal operations and by remotely moving the sources to positions nearby the detectors only during calibration runs.
@@
SABRE South uses motorised calibration deployment systems for the calibration of both the crystal and the liquid scintillator detectors, denoted RCS1 and RCS2, respectively (Fig. 17).
Each system is composed of a steel mount, a stepper motor, a calibration probe suspended through a wire, an alignment reel, and sensors for position measurement. The stepper motor is used to wind and unwind a wire around a spool, thus moving the calibration probe up and down. The wire passes around a centring reel, so that the probe can accurately slide into an aluminium pipe, which guides it to the calibration position. The systems are controlled by the online slow control system using the EtherCAT protocol and are capable of moving the radioactive sources between the calibration position and the normal operation position in minutes. No human intervention is needed in the experiment during operation. Access will only be required in the unexpected case of parts failure.
@@
Figure 17: Close-up views of the radioisotope calibration systems for (left) the crystal detectors and (right) the veto system.
@@
Three units of each RCS are installed on top of the veto vessel, shown in Fig. 13, so that the detectors are evenly irradiated and the PMT signals are comparable. Each triplet of RCS units is equidistant from the axis of the vessel and 120 degrees apart. The RCS1 units are positioned so that each calibration source is equidistant from a triplet of crystal detectors when lowered into the calibration pipe (150 mm from the crystal axis).
@@
The RCS2 units are mounted on top of the three outer flanges (151.6 mm diameter) at a distance of 800 mm from the centre of the vessel. Under consideration is the installation of a steel cylinder (151.6 mm external diameter, 16.8 mm internal diameter and 100 mm height) on top of the flange to house the calibration sources and shield its radiation during normal data acquisition. During calibration, the sources of each RCS2 system are lowered to the vessel mid-plane (the same height as the centre position of the crystals) at a position equidistant from pairs of veto PMTs. Each veto PMT is at most 1.5 m from any radiation source.
The RCS2 calibration probes enter the vessel through aluminium pipes in order to keep the LAB sealed and separated from the calibration probe and the external atmosphere.
The sources used for the RCS1 system will be 109Cd, 210Pb, and 241Am, while RCS2 will use 137Cs, 133Ba, 22Na, and 207Bi. The simulated energy depositions (after smearing) for each system is shown in Fig. 18 (for RCS1, this applies only to a single crystal). The muon veto is calibrated with a 60Co source.
Figure 18: Left: Energy depositions for the sources used to calibrate the LS detector folded in with the detector resolution. Right: Energy depositions for the sources used to calibrate the crystal detectors, folding in detector resolution.

### v3

@@ 4.2 R11065 PMTs
The R11065 PMTs were chosen for their high sensitivity to single photons, low noise rates, and low radioactive background. Only PMTs with quantum efficiency (QE) $>30\%$ were procured, so to achieve a high single-photon detection efficiency. Custom PMT voltage dividers are used to minimise their contribution to the background. These bases will follow the resistance ratios recommended by Hamamatsu for this type of PMT. SABRE South has decided on a design that will use a positively biased voltage divider. The circuit diagram and an image of the PCB for the voltage divider design are shown in Fig. 8. The final design is expected to use small ceramic capacitors. The nominal design does not include an onboard or offboard preamplifier. Preamplifiers are being considered for a future upgrade, but would not be installed onboard.
Figure 8: Left: The circuit diagram for the final R11065 base design. Right: One of the PCBs to be used for the base circuit, which will be attached to an R11065 PMT.
The pre-calibration program for the R11065 (crystal detector) PMTs focuses on confirming their single-photon detection efficiency, characterising their single-photo electron (PE) response, and characterising the noise as a function of temperature and voltage. These parameters are used as an early calibration input to SABRE South data analysis and as an input into the digitisation stage of detector simulations. The requirements for a quantum efficiency larger than $30\%$ (at a wavelength of 420 nm) and excellent signal-to-background discrimination for single photons are driven by the light yield requirements (of $\sim$ 12–14 PE/keV) on the crystal detectors to achieve a threshold of 1–2 keV. Several essential measurements will be completed for all PMTs under strict cleanliness protocols. These are: relative quantum efficiency, gain, dark rate, and afterpulse rate at nominal voltage. Further characterisation of single-photon response and noise, including temperature and bias voltage dependence, is performed on a limited set of PMTs to limit overall exposure. As SABRE operates at room temperature, dark counts due to thermionic emission and temperature dependence of this dark rate are essential factors that are more easily measured in a dedicated setup than in-situ. The pre-calibration datasets allow for the development of data analysis techniques for discrimination of thermionic emission from low-energy events in the crystals, near the 1–2 keV region. There have previously been reports [41, 42, 43] of light emission likely due to anode and dynode glow following a sufficiently intense light signal in the R11065 and the related R11410 PMT types. The R11410 and R11065 PMTs differ only in the choice of photocathode material. Such light emission could introduce an additional background in SABRE and therefore needs to be characterised if present.
@@ 4.3 Cleanliness protocols
All components of the crystal enclosures are cleaned and vacuum baked prior to final assembly to ensure that there is no moisture remaining on the components. The crystal PMTs, cables, and bases are cleaned with lint-free isopropyl alcohol (IPA) soaked cleanroom wipes and a high pressure nitrogen gun to remove any debris. To prevent radon plate-out, all crystal detector module components are protected in two-layered radon bags for storage between cleaning and their installation in the final detector modules. The assembly of the base/cables and their mounting on the PMTs is performed in an ISO6 cleanroom environment. Assembly of the crystal modules is performed in a nitrogen flushed glove box.
@@ 4.5 Gas Handling System
Figure 11: The exhaust system gas panel undergoing laboratory testing at the ANU.
@@ 5 Liquid scintillator veto
SABRE South’s active veto system serves to identify background events originating from radioactive decays within the crystals that may mimic a signal and to act as passive shielding for background external to the crystals. For example, 40K, a key impurity, decays through the emission of an Auger electron with an energy in the range of 1–6 keV coincident with a 1.5 MeV $\gamma$ ray. This decay process proceeds via K-shell electron capture, by which the 40K becomes an excited state of 40Ar and emits a 3 keV X-ray and an Auger electron, followed by the decay of the excited state and the emission of a 1.5 MeV $\gamma$ ray. Previous experiments have accounted for this by employing double-coincidence rejection, where it was assumed that the escaping gamma ray would be detected by another crystal in the sensitive mass [44]. However, this detection mechanism is not fully efficient, as demonstrated by the peak at around 3 keV in the DAMA data where events from 40K are expected [45].
By immersing the NaI(Tl) crystal detector modules in a liquid scintillator (Fig. 12), high-energy $\gamma$ rays produced by radioactive decays can be detected and the decay event vetoed. The SABRE South full background simulation demonstrated that the veto can reduce background from important 40K contamination by a factor of 10 in the region of interest and suppress the total experimental background by 27% [11]. While the 210Pb contamination has a large contribution to the total background, it is not possible to veto as $\alpha$ particles are predominantly emitted along the decay chain, as well as low energy $\gamma$s that do not reach the liquid scintillator.
The trigger logic for the veto is being designed to optimise veto efficacy with minimal deadtime, with information such as multi-PMT coincidence as well as charge-based and time-based information.
@@
SABRE South uses approximately 11,600 L of liquid scintillator that is a mixture of linear alkylbenzene (LAB), sourced from the supplier to the JUNO experiment, doped with approximately 3.5 g/L of the fluorophore PPO (2,5-diphenyloxazole) and 15 mg/L of bis-MSB (1,4-bis(2-methylstyryl) benzene). The fluorophore concentrations are subject to further optimisation to ensure low absorption [46]. Purification of the liquid scintillator is not planned during operation. The veto detector uses Hamamatsu 204 mm R5912 PMTs with oil-proof potted bases. The nominal design uses a baseline of 18 PMTs, but it is possible a total of 32 could be installed with the inclusion of PMTs from the decommissioned Daya Bay experiment. The PPO and Bis-MSB are required to mitigate the self-absorption properties of LAB [46] and maximise the total light yield. The wavelength of light output of LAB alone is around 300 nm [47]. In the presence of fluorophores the maximum wavelength of emission is 422 nm, which is closer to the peak quantum efficiency of the PMTs. The key components of the liquid scintillator veto detector are summarised in Table 4.
@@
The liquid scintillator is effective in rejecting events in which electrons, neutrons, and photons deposit more than 20 keV of energy in the scintillator. For the baseline scenario with 18 veto PMTs, this enables the detection of energy depositions with 50% efficiency at a threshold of 20 keV, rising to 99% at 80 keV, based on simulations of energy depositions from crystal background sources in the veto. This expected performance is important if the DAMA/LIBRA signal is to be excluded, as it provides a means of understanding what background processes might cause a DAMA/LIBRA-like modulation through the use of techniques such as pulse-shape discrimination. This has been tested conceptually using a liquid scintillator prototype detector, containing approximately 40 L of LAB-based liquid scintillator and a single R5912 PMT, in which neutron and $\gamma$ ray signals could be discriminated. This suggests that the same approach could be applied to the full liquid scintillator veto, particularly since the arrival of 16 additional PMTs from the decommissioned Daya Bay experiment gives us the opportunity to increase the number of PMTs in the detector from 18 to 32, thereby increasing photosensor coverage and light yield/light collection efficiency. Given a 0.75 PE/keV position averaged detection probability for detection by any given PMT, an energy threshold for pulse shape discrimination can be estimated to be as low as 150–200 keV. This depends on the method of pulse shape discrimination, with advanced multivariate machine learning techniques likely to reduce the threshold. Position reconstruction in the liquid scintillator veto is also a goal for the experiment and is being explored. This is useful for identifying events in the veto that may have originated from the crystals, thereby aiding the veto process via fiducialisation of a volume around the crystals.
@@ 5.1 Vessel design
The vessel structure can be divided into three sections: the central barrel and the two torispherical caps located at the top and bottom of the vessel. To ensure that the vessel is small enough to be safely transported to SUPL on the back of a vehicle and large enough for MeV-scale photons to be fully absorbed inside the liquid scintillator, an outer diameter of 2600 mm (at the middle of the barrel) and a total height of 3322 mm (inclusive of the top flange) were chosen. A diagram of the vessel is shown in Fig. 13. Rows of stud bolts are welded to the internal wall of the vessel to mount the PMTs and reflector. A central flange (“top flange”, see Fig. 14), is located at the top of the vessel, which is flanked by a circle of smaller side flanges. The central flanges are used for the insertion of the crystal enclosures, fluid handling, cable feedthroughs, and calibration systems.
@@ 5.2 Assembly
Figure 15: Engineering models of the central and lateral mounting units. The mounts are nominally made of aluminium. Labels 1–3 highlight the central mount frame; labels 4 and 5 point out the bulb and oilproof base of the R5912 respectively; and labels 6–8 indicate the clamps in which the PMT is mounted. The mounts will be fastened to two stud bolts, at the top and bottom of each mounting bracket, whilst the clamps will be fastened with stainless steel nuts and bolts.
@@ 5.3 Cleanliness protocols
The vessel must be pickled internally because an oxide layer is expected to have formed between manufacturing and assembly, which may contain radioactivity and may not be compatible with the liquid scintillator. Pickling is performed using a corrosive solution with a concentration that is balanced enough to corrode the oxide layer and not the steel itself. Following this, the vessel is rinsed multiple times with a solution of de-ionised water and Alconox detergent, followed by a rinse with de-ionised water. Rinsing ceases once the outgoing water/solution reaches a required resistivity of $\geq 14$ MOhm cm. The R5912 veto PMTs and Lumirror, along with all other internal components, follow a similar cleaning procedure to the R11065 PMTs, which is also performed in a cleanroom environment. The installation of internal components is performed in a clean room-like environment with radon reduction, with all installation personnel wearing clean-suits, boot-covers, gloves, and hairnets. Pumping and transfer of the liquid scintillator into the vessel will be performed clean, with the liquid scintillator sparged with HPN once the vessel has been filled. The liquid scintillator was transported and stored in a new ISO tank made of 316L steel, and blanketed in nitrogen. The transferral and cleaning of this tank was handled by the liquid scintillator producer, which was direct from the distillation port and acceptably clean. In case of contamination of the liquid scintillator during filling or transfer we will pause the filling process. There is a total of 17 kL of liquid scintillator in the ISO tank, and since only 11.6 kL is required for the LS veto. The remainder can be used for liquid transfer commissioning.
@@ 5.4 R5912 PMTs
Owing to the sparse array of PMTs relative to the volume of liquid scintillator, the probability of photoelectron (PE) detection is 0.2–0.3 PE/keV/PMT. Consequently, the pre-calibration of these PMTs is driven by the need to be sensitive to low numbers of PEs, to inform our simulation/digitisation and thus analysis, and also by the ability to deploy a degree of position reconstruction and particle ID in the liquid scintillator veto. With such reconstruction capabilities, we can disentangle our background contributions and pursue other physics using the veto sub-detector — for example, studies of neutron spallation induced by muon tracks, or potential detection of supernova neutrinos (in which tagging the neutron of an inverse beta decay could improve our detection efficiency). There is no magnetic shielding surrounding these PMTs.
@@ 6 Muon veto
Figure 17: Muon detectors in SUPL, assembled in ‘telescope’ mode, which is the configuration required to measure the muon flux and angular spectrum (from two directions).
The digitiser used for the readout is a CAEN V1743, which was specifically chosen for its high sample rate of 3.2×$\times 10^{9}$ samples/s (which complements the TTS of the PMTs) and resolution of $<$ 8 ps RMS [50]. The timing resolution of this system allows for accurate position reconstruction along the length of each muon paddle, where the time difference between signals in each PMT is used to determine position. Measurements performed on the surface have found that this can be performed at a position resolution of approximately 5 cm. The efficiencies of each detector have also been measured on the surface to be at least 99% for muons that fully traverse the thickness of the detector. The surface flux and statistical uncertainty at the University of Melbourne was measured to be 131.3 $\pm$ 0.4 s-1m-2. At SUPL we expect the muon rate for each panel (1.2 m2), with statistical uncertainties, to be 37.8 $\pm$ 4.3 d-1, and 302.7 $\pm$ 34.0 d-1 for all eight panels with an overall coverage of 9.6 m2, based on a measurement with a smaller detector. Figure 18 shows the expected energy and angular distributions for muons reaching SUPL compared to other underground labs, where the distributions are based on calculations from models provided in Ref. [51], and the points for SUPL are derived from simulations using Geant4 [52, 26] and CRY [53]. Muons will always be vetoed, and there should be no muon background in the crystals. Highly angular muons may miss the muon veto, but will still be vetoed by the liquid scintillator. Likewise, if vertical muons miss the liquid scintillator by passing through conduits/other detector material, the vertical coverage provided by the muon detectors will provide the veto.
Figure 18: Left: Simulation results showing the normalised angular spectrum of muons that reach SUPL, which we aim to experimentally verify with the muon detectors assembled in ‘telescope’ mode in SUPL. One could use angle projected by the size of the muon veto and this distribution to estimate the number of muons that will hit the liquid scintillator, and miss the muon detectors. These muons will be vetoed by the liquid scintillator instead. Right: The simulated energy distribution of muons that reach SUPL. This can be expected to be the energy range of muons to be tagged by the muon veto, although the muon detectors will not be used as a spectrometer.
@@ 7 Calibration systems
The SABRE South sub-detectors require periodic calibration, which is particularly important for accurate time-dependent measurements. Calibration refers to the process of bias correction and verification of the accuracy of a detector’s measurements. The liquid scintillator veto and NaI(Tl) calibration systems for SABRE South are designed for remote operation, and are located within the shielding structure. The response of the detector to particle interactions can change with time, due to: (i) environmental factors such as temperature, humidity, mechanical stress, and vibrations (monitored by the slow control system outlined in Sec. 8.5); (ii) electrical changes, including variations in the supply voltage and electromagnetic interference; (iii) detector ageing, where over time the optical properties of scintillators can deteriorate (e.g. reduced absorption length or light yield), and the PMTs can experience changes in gain, detection efficiency and other features used for pulse shape discrimination.
SABRE South uses radioactive sources to calibrate the NaI(Tl) detectors, the liquid scintillator veto, and the muon detector. Additionally, the liquid scintillator veto is calibrated using an optical (pulsed laser light) system. The two calibration systems for the liquid scintillator veto are complementary: the radioactive calibration system provides accurate high-energy calibration lines, while the optical system is capable of sending low-intensity, high-frequency pulses to test the single photoelectron and timing response of the PMTs. Calibration with radioactive sources requires data runs of a few hours; other NaI(Tl) experiments typically have 3–4 hour calibration runs every 2 weeks [54]. The laser calibration can be conducted more frequently in short bursts, as it does not require any mechanical adjustments to the system. The frequency of calibration for both systems will be adjusted as needed to optimise detector stability and live-time.
@@ 7.1 Radioisotope calibration system
To calibrate the detectors, radioactive sources must be placed close enough to produce a signal distinguishable from other detected signals during calibration runs. However, during normal data acquisition, the signal from the calibration source must be kept negligible. This is managed by keeping the radioactive sources shielded and at a distance from the detectors during normal operations, then moving them closer to the detectors only during calibration runs. Simulations show that by using sources with sufficiently low activity and ensuring adequate shielding, the dead time of the veto is minimally affected. This approach eliminates the need to disassemble the shielding for each calibration run, maximising detector live-time while maintaining efficient calibration.
@@
SABRE South employs motorised calibration deployment systems for the precise calibration of both the crystal detectors and the liquid scintillator detectors, referred to as RCS1 and RCS2, respectively (Fig. 19).
Each radioactive calibration system consists of a steel mount, a stepper motor, a calibration probe suspended by a wire, an alignment reel, and position measurement sensors. The stepper motor controls the winding and unwinding of the wire around a spool, enabling the precise vertical movement of the calibration probe. The wire is guided through a centring reel, ensuring the probe slides accurately into an aluminium pipe that directs it to the calibration position.
The systems are operated via the slow control system using the EtherCAT protocol, which allows radioactive sources to move between the calibration and normal operation positions in minutes. This process is fully automated. Access to the system will only be necessary in rare cases of equipment failure.
@@
Figure 19: Close-up views of the radioisotope calibration systems for (left) the crystal detectors and (right) the liquid scintillator veto system.
@@
Three units of each motorised calibration deployment system are installed on top of the veto vessel, as shown in Fig. 14, so that the detectors are evenly irradiated and the PMT signals are comparable. Each triplet of RCS units is equidistant from the axis of the vessel and 120 degrees apart. The RCS1 units are positioned so that each calibration source is equidistant from a triplet of crystal detectors when lowered into the calibration pipe (150 mm from the crystal axis).
@@
The RCS2 units are mounted on top of the three outer flanges (151.6 mm diameter) at a distance of 800 mm from the centre of the vessel. Under consideration is the installation of a steel cylinder (151.6 mm external diameter, 16.8 mm internal diameter and 100 mm height) on top of the flange to house the calibration sources and shield their radiation during normal data acquisition. During the calibration runs, the sources of each RCS2 system are lowered to the vessel mid-plane (the same height as the centre of the crystals) at a position equidistant from pairs of veto PMTs. Each veto PMT is at most 1.5 m from any radiation source. The RCS2 calibration probes are designed to enter the vessel through aluminium pipes, ensuring that the LAB remains sealed and isolated from both the calibration source and the external atmosphere.
The RCS1 system uses the following radioactive sources 109Cd and 241Am, each with an activity of approximately 10 kBq. This activity level is chosen because the background radiation is small compared to other sources of background. For RCS2, the system uses higher-energy sources, 137Cs, 133Ba, 22Na, and 207Bi. These sources produce more significant background signals, so their activity is kept below 1 kBq to minimise this effect. The simulated energy depositions for each system are shown in Fig. 20, which illustrates the energy depositions for the RCS1 system using a single crystal. The muon veto is calibrated with a 60Co source.
Figure 20: The left figure shows the energy depositions for the sources used to calibrate the LS detector, with the detector resolution taken into account. The right figure shows the energy depositions for the sources used to calibrate the crystal detectors, also considering the detector resolution. Both sets of spectra have been smeared to match the expected resolution of each detector and are zero-suppressed, with error bars representing the statistical uncertainty. The distributions are normalized to reflect the number of disintegrations expected in one hour for 10 kBq sources. The inclusion of 210Pb was a test to evaluate its viability as a potential source.